OpenStudy (anonymous):

How do I simplify y^7/y^5

OpenStudy (anonymous):

Still here

OpenStudy (anonymous):

\[y^7/y^5 = y^7-y^5 => y^2\]

OpenStudy (john_es):

\[y^7/y^5=y^{7-5}=y^2\]

OpenStudy (anonymous):

Yes sorry about that, 5 am lol it's y^7-5, nice catch John.

OpenStudy (anonymous):

Thank you